OpenStudy (anonymous):

-7/8 (-2,6) equation of a line with slope y=mx+b

OpenStudy (anonymous):

Is -7/8 the slope ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok do you know the formula of creating a line by knowing the slope and a point ?

OpenStudy (anonymous):

i can do these as long as its not a fraction one i cant grasp the fraction and at the end on how to know whether to add or subtract

OpenStudy (anonymous):

i know the end will come out to be y=-7/8x and the rest i have no clue

OpenStudy (anonymous):

ok simply you have to substitute by m= -7/8 in the equation y=mx+b

OpenStudy (anonymous):

i understand that. -7/8 (-2)+b, from here im confused

OpenStudy (anonymous):

\[y=\frac{ -7 }{ 8 }x+b\]

OpenStudy (anonymous):

then to find b let x=-2 and y=6

OpenStudy (anonymous):

ok got that

OpenStudy (anonymous):

got b ?

OpenStudy (anonymous):

no after that last step is where i get confused

OpenStudy (anonymous):

let x=-2 and y=6 in the equation above and calculate b

OpenStudy (anonymous):

14/8 ?

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

see this is where i have no clue how to do this

OpenStudy (anonymous):

i multiplied -7/8 (-2)

OpenStudy (anonymous):

\[6=\frac{ -7 }{ 8 }*-2+b\]

OpenStudy (anonymous):

calculate b from this equation

OpenStudy (anonymous):

i got 14/8

OpenStudy (anonymous):

wrong b=34/8

OpenStudy (anonymous):

how

OpenStudy (anonymous):

the calculator gets 17/4

OpenStudy (anonymous):

it is 6-14/8=34/8